# Augmented Lagrangian Solvers for Poroelasticity with Fracture Contact Mechanics

### v1

@@ 3. Convergence check:
  if $(\bm{r}^{(k+1)},\bm{\lambda}^{\text{trial}})$ fulfills convergence criterion then
@@
   Define the new friction bound $b^{(k+1)}=-F\lambda_{n}^{(k+1)}$.
@@
   Set new guess to $(\bm{r}^{(k+1)},\bm{\lambda}^{(k+1)})$, set $k=k+1$ and go to 2.
Remark 1 (Derivative of max-function).
We have observed empirically that the return map will quite frequently result in cases where $\lVert\bm{\lambda}_{\tau}^{(k+1)}+c[[\dot{\bm{u}}^{(k+1)}]]_{\tau}\rVert=b^{(k+1)}$, i.e. the two arguments of the maximum-function of (24) are equal. In this case, GNM-RM was found to be more robust if the Jacobian corresponding to the second argument is chosen, and so this is done in all of the simulations in section 4. For GNM and IRM, neither of the two options showed a significant increase or decrease in the robustness of the solvers. Therefore, we arbitrarily choose the Jacobian corresponding to the second argument in case of equal arguments also for GNM and IRM.
It should be noted that there is a fundamental difference in the purpose of the return map for GNM-RM, compared to IRM. For IRM, the return map (31) is used in place of the complementarity functions (23) and (24), while for GNM-RM, the same return map is used on top of solving the complementarity functions (23) and (24). Hence, for IRM the return map is an essential part of the solution strategy for finding tractions and displacements that fulfill the contact conditions. However, this is not the case with GNM-RM, as fulfillment of the contact conditions is already handled by performing Newton iterations on the complementarity functions (23) and (24), and so the return map for GNM-RM functions merely as stabilization of GNM.
@@ 3.4 Finite volume discretization
We conclude this section with a brief summary of the spatial discretizations used in this work, and the corresponding practical implementation of algorithms 1-3. We follow the discretization schemes described in [8] and the references therein. In short, we employ a family of multi-point finite volume approximations for diffusive terms, while an upwinding scheme is used for advective terms. The grids are simplicial and conforming to the fractures, in the sense that lower-dimensional cells coincide with higher-dimensional faces. In this fully discrete formulation, the residual system (22) becomes a nonlinear system of algebraic equations, while the complementarity functions are enforced cell-wise, resulting in another system of nonlinear algebraic equations. The implementation of the generalized Newton method on the full system (or on the regularized systems in the case of IRM) is now straightforward. In GNM-RM, the return maps are performed cell-wise.
@@ 4 Numerical experiments
In this section, we will compare the performance of the three nonlinear solvers presented in section 3, GNM, IRM and GNM-RM, through a series of two- and three-dimensional numerical simulations. Specifically, we simulate pressurized fluid injection in a fractured reservoir, mimicking hydraulic stimulation of a fractured geothermal reservoir in crystalline rock. All of the simulations are run using the mixed-dimensional simulation framework PorePy [30]. Runscripts for reproducing the results of this section can be found in the Docker image in [39]. The precise vertex coordinates of the fractures used in the simulations are reported in the supplementary material of this article.
A central focus of the experiments is to test the sensitivity of the solvers to the value of the augmentation parameter $c$. There are some recommendations on how to choose this parameter from the literature on purely mechanical contact problems. For IRM, it is well established that higher $c$-values lead to faster convergence of the outer loop, but usually with more iterations per inner Newton loop [24, 16]. Hence, there is often an optimal intermediate $c$-value that gives a compromise between the two effects. For GNM, it is recommended in [27, 40] to choose the parameter such that the two terms involving $\bm{u}]]$ and $\bm{\lambda}$ in the maximum-functions of (23) and (24) are balanced. However, these recommendations may not be easy to implement in practice. Firstly, it is difficult to assess a priori the optimal intermediate $c$-value of IRM, as shown in [24]. Secondly, it may be difficult to estimate the ratio of $\bm{u}]]$ and $\bm{\lambda}$. For simple fracture network geometries (e.g., isolated fractures), this ratio can be estimated from the boundary conditions and material parameters. However, fracture contact mechanics can alter the stress field such that, for intersecting or close fractures, the actual ratio is far from such estimates. The discrepancies may increase further when adding poromechanical effects. These difficulties motivate the testing of a large range of $c$-values, and assessing the robustness of the solvers in terms of their sensitivity to the value of $c$.
For each of the solvers, we report on the total number of nonlinear iterations needed for convergence. By this we mean the total number of linear systems that was solved for the duration of the simulation, keeping in mind that for IRM we additionally need to sum over all outer iterations of the nested loop. For the convergence criteria, we require the Euclidean norms, scaled by cell volumes, of both the nonlinear iteration increments and residuals to be below $10^{-8}$. For IRM, the same convergence check is done both in the outer and inner loops (recall that the regularized system of the inner loop is solved by GNM). We set the maximal number of allowed iterations per nonlinear system as 100, and simulations where this threshold is exceeded are declared as non-convergent and labeled by ”NC”. Another way in which the solvers can fail is if the iterations diverge to infinity. We consider a simulation as diverging to infinity if the norm of the residual exceeds $10^{5}$ and label it by ”Div”. Finally, we set the maximum number of allowed iterations of the outer loop of IRM to be 150, and simulations exceeding this threshold is labeled by ”NCO” (non-convergent outer). If an IRM simulation is labeled ”NC” or ”Div”, it means that GNM in the inner loop failed to converge or diverged to infinity.
An appropriate scaling of the units is used to improve the conditioning of the nonlinear systems. It is found empirically that scaling the unit of mass by $10^{9}$ kg results in the primary variables being close to the same order of magnitude, and generally improves the robustness of all solvers. Hence, this unit scaling is used in all simulations. The material constants used in the simulations are listed in table 2.
@@
Residual aperture $a_{\mathrm{ref}}$ $5.0\cdot 10^{-4}$ m
Dilation angle $\psi$ 3, 5 or 7 degrees
@@ 4.1 Two-dimensional simulations of fluid injection
In our first series of test cases, we restrict ourselves to two dimensions, and neglect gravitational effects. A domain of size $2000\times 1000$ m is considered, with the fracture network as shown in figure 2. The centermost cell of one of the fractures will be kept at a constant higher pressure, which can be thought of as simulating an injection well (marked by a white dot in the figure). We consider an anisotropic stress regime where pressurization of the injection well will cause fractures to slip and open, creating complex dynamics that challenges the nonlinear solvers.
@@
For the flow boundary conditions, we impose pressure values of 20MPa on all external boundaries. For the mechanical boundary conditions, we fix the bottom boundary (zero displacement), while compressive tractions are imposed on the remaining boundary sides. The traction values correspond to a stress tensor with non-zero values $\sigma_{xx}=-30$ MPa and $\sigma_{yy}=-50$ MPa. These pressures and anisotropic stress values are representative for crystalline rocks around 2 km depth. For all simulations, suitable initial values of the variables are found by first imposing a constant pressure field of 20 MPa everywhere, and solving only the purely mechanical subproblem. The solution to this simpler problem is used as the initial guess for the nonlinear solvers.
We test the effect of imposing four different values of the pressure in the injection well: 20.01, 21, 28 and 30 MPa. For the highest injection pressure, we also test the effect of varying the dilation angle $\psi$. In all cases, we use a grid containing 33830 total cells (this includes matrix, fracture and interface cells). All simulations are run for a single time step of 0.1 days. We consider only a single time step, as the first time step was typically more difficult than subsequent time steps. Figure 3 shows the performance of the solvers for different values of the augmentation parameter $c$. Note that the parameter is given in units of GPa/m, in accordance with our scaling of the unit of mass. A detailed plot of the contact states and pressure profile for the 30 MPa case with $\psi=5^{\circ}$ is provided in figure 2, showing how the elevated pressure in the injection well causes some of the fractures to slip and open.
Figure 2: Visualization of the simulation with an injection pressure of 30 MPa and a dilation angle of $5^{\circ}$, showing the fracture network used in all simulations of section 4.1, and the position of the injection well (the latter is marked by the white dot). Left: Contact states (as defined by (28)) after the initialization stage, but before the injection well is activated. Here, fracture slip is measured relative to a configuration with no deformation. Right: Contact states and pressure profile after activating the injection well and taking a time step of 0.1 days. Here, fracture slip is measured relative to the state after the initialization, i.e. the left figure
Figure 3: Number of nonlinear iterations for one time step for the example of section 4.1 with different injection pressures (subfigures 3(a)-3(d)) and for different dilation angles (subfigures 3(d)-3(f)). NC and Div denote simulations in which the solver did not converge within 100 iterations or diverged to infinity, respectively (for IRM, NC and Div have the same meaning, but applied to GNM in the inner loop). NCO denotes IRM simulations where the outer loop exceeded 150 iterations
From figure 3 it is clear that increasing the injection pressure adds to the difficulty of the problem, with an increase in both the iteration counts, as well as the number of non-convergent cases. For IRM we also have the problem of the outer loop being significantly slower for lower values of $c$, with some cases terminating due to the outer loop using more than 150 iterations. In the simplest case, figure 3(a), corresponding to an injection pressure of 20.01 MPa, we can observe the optimal value of $c=10^{3}$ occurring as a compromise between faster outer loops but more difficult inner Newton loops, with the inner Newton loop simply failing to converge for the highest c-value. As previously mentioned, such an optimal $c$-value is well established in the contact mechanics literature. However, the pattern is immediately broken in the next simplest case, figure 3(b) (injection pressure 21 MPa), as the inner Newton loop of IRM fails to converge in all cases, except when $c=1$. For the more difficult cases, the inner Newton loop does not converge at all. The fact that IRM converges in far fewer cases than GNM suggests that regularizing the complementarity functions actually increases the difficulty of the system. It is clear that IRM is outclassed by the other solvers here. Therefore, we will omit IRM from the rest of the numerical experiments.
GNM appears to handle the smaller values of $c$ better than the larger ones, with $c$-values of 10 and above only being successful in the simplest case with the 20.01 MPa injection pressure. For the three lowest injection pressures, the method is quite consistent, in terms of the lower $c$-values consistently resulting in convergence, and the number of iterations being relatively similar across these values. However, the method becomes more unpredictable for the most difficult cases with the 30 MPa injection pressure, both in terms of which $c$-values that result in convergence, and the number of iterations needed. The method overall seems to have the highest chance of converging for $c$-values from around $10^{-2}$ to $10^{-1}$ GPa/m, as these are the values with the overall fewest number of non-convergent cases. A closer inspection of the solution values reveals that the ratio $\bm{\lambda}/[[\bm{u}]]$, when both are nonzero, is typically in the range from $10^{-2}$ to $1$ GPa/m for the examples of this section. Hence, the recommendation from [27, 40] of choosing $c$-values that balance the scales of $\bm{u}]]$ and $\bm{\lambda}$ is a relatively good fit to these simulations. However, as mentioned, convergence is far from guaranteed for the most difficult cases, even for $c$-values in this range.
Moreover, figures 3(e) and 3(f) show that GNM can also be sensitive to small changes in the input data, as slight variations in the dilation angle can cause a convergent case to stop converging, or vice versa. It appears that decreasing the dilation angle increases the difficulty of the problem in this case, but the opposite effect has also been reported for other setups [41]. Finally, GNM-RM has the overall best performance out of the three solvers, converging across the largest range of $c$-values. In particular, it handles the larger $c$-values much better than the other two solvers. However, like GNM, it also becomes unpredictable for the most difficult setups, in terms of which $c$-values that result in convergence, and the number of iterations needed. The performance of GNM-RM can also be sensitive to small changes in the input data, such as the dilation angle, as seen in figures 3(e) and 3(f).
A common problem with Newton methods is cycling; in several of the cases in figure 3 where GNM and IRM fail to converge, but without the residual diverging to infinity, the problem is cycling. It appears to be particularly prevalent when a high $c$-value is used. An example of this is shown in figures 4 and 5, in which data from the simulation with an injection pressure of 28 MPa and a $c$-value of $10^{2}$ is plotted, corresponding to the results of the seventh column of figure 3(b). Figure 4(a) plots the norm of the total residual vector for the three solvers. The cycling of GNM and IRM is illustrated by repetitive patterns in the residual history. In figure 4(b), we show the effect of switching from GNM to GNM-RM, by turning on the return map after 20 iterations. Switching to GNM-RM results in convergence, indicating that GNM-RM can successfully avoid the cycle.
Moreover, the cycling of GNM and IRM occurs nearly always between different contact states, as defined by (28) for GNM and by (35) for IRM. This corresponds to cycling between different arguments of the maximum functions in the complementarity functions. Figure 5 plots the total number of fracture cells residing in each contact state for the three solvers. We observe that for both GNM and IRM, some of the cells cycle between the different states, while GNM-RM avoids the cycle and converges on a fixed number of cells in each contact state.
Figure 4: Residual plots for the different solvers for the simulation with an injection pressure of 28 MPa and $c$-value of $10^{2}$, corresponding to column 7 of figure 3(c). Left: Residual norms at every iteration for GNM, IRM and GNM-RM. Only the first 50 iterations of GNM and IRM are shown. Right: The same GNM cycle as in the left figure, showing how the cycle can be broken by switching to GNM-RM
Figure 5: Number of cells in each contact state for the different solvers for the simulation with an injection pressure of 28 MPa and $c$-value of $10^{2}$, corresponding to column 7 of figure 3(c). The contact states are defined by the sets (28) for GNM and GNM-RM, while for IRM, the contact states are ”regularized” and defined by (35)
Next, we check the scalability of GNM and GNM-RM with respect to the grid size. Using the same fracture network as before, a sequence of four grids is considered: Starting from the grid with 33830 total cells used in the previous simulations, we consider finer grids of 64068, 121501 and 199282 total cells, respectively. The grids are constructed independently of each other. Special attention must be paid to the injection well in order to make the physical setup as similar as possible for the different grids. For each grid we will consider a fracture cell as an injection cell if its center is within the bounds of the injection cell of the coarsest grid, thus ensuring the spatial extent of the elevated pressure corresponding to the injection to be close to constant for all grids.
We run the grid refinement for three cases: One simpler case with an injection pressure of 21 MPa and two more difficult cases with injection pressures of 30 MPa. In the simpler case, we set $c=10^{-2}$ and $\psi=5^{\circ}$, corresponding to the third column of figure 3(b). For the other two cases, we use $c=10^{-2},\psi=5^{\circ}$ and $c=10^{-1},\psi=3^{\circ}$, corresponding to the third and fourth columns of figures 3(d) and 3(e), respectively. To make the grid dependency clearer, we now run more than one time step; specifically, we run the simulations for ten time steps, using a constant step size of 0.1 days. The cumulative iteration counts across all time steps are reported in figure 6.
Figure 6: Cumulative number of iterations over ten time steps, as a function of the grid size.
The results of the simplest case (21 MPa injection pressure) indicate a weak dependency on the grid resolution, showing a modest increase in the total number of iterations for finer grids for both GNM and GNM-RM. A similar weak grid resolution dependency was reported for GNM in [24] for purely mechanical contact problems in non-fractured media. The more difficult cases (30 MPa injection pressure) deviate from this pattern, however, including some non-convergent cases. Note that all of the non-convergent cases fail to converge during the first time step. These observations are similar to the pattern observed in figure 3, namely that as the problem becomes more complex (in this case meaning higher injection pressure), the nonlinear solvers become more unpredictable, and deviate from established results in the contact mechanics literature. Moreover, the results of figure 6 show that changes to the grid structure, caused by the grids being constructed independently, can change the convergence behavior. The influence of the precise grid structure on the convergence behavior for problems of fractured porous media has been noted before in [42].

### v2

@@ 3. Convergence check:
  if $\lVert\bm{x}^{\text{trial}}-\bm{x}^{(k)}\rVert<\text{TOL1}$ and $\lVert F(\bm{x}^{\text{trial}})\rVert<\text{TOL2}$ then
@@
   Define the new friction bound $b^{(k+1)}=-\mu\lambda_{n}^{(k+1)}$.
@@
   Update guess to $\bm{x}^{(k+1)}=(\bm{r}^{(k+1)},\bm{\lambda}^{(k+1)})$, set $k=k+1$ and go to 2.
We note that the return map in GNM-RM is explicitly computed after each generalized Newton iteration, which is in contrast to IRM, where the return map is implicitly defined through the regularized complementarity functions. GNM-RM is hence distinctly different from an Uzawa method like IRM, and should instead be interpreted as a quasi-Newton method.
GNM-RM is conceptually similar to the so-called projected Newton method [projected_newton] from the optimization literature. This method solves optimization problems with inequality constraints by following each Newton iteration with a return map (projection) onto the feasible set. However, the return map plays a conceptually different role in GNM-RM, compared to both the projected Newton method of [projected_newton], as well as IRM. For the latter two methods, the return map (projection) is performed in order to ensure the fulfillment of the inequality constraints. For GNM-RM, on the other hand, the fulfillment of the inequality constraints, i.e the contact conditions, is already handled by performing generalized Newton iterations on the complementarity functions (26) and (27). Hence, the return map of GNM-RM functions merely as stabilization of GNM.
@@ 3.4 Finite volume discretization
We conclude this section with a brief summary of the spatial discretizations used in this work, and the corresponding practical implementation of algorithms 1-3. We follow the discretization schemes described in [stefansson2023flexible] and the references therein. In short, we discretize the conservation laws by a family of cell-centered finite volume schemes. For the fluid mass balance equation, the standard multi-point flux approximation is used, while for the momentum balance, we use the multi-point stress approximation, first introduced in [nordbotten2014cell]. Advective terms are discretized by an upwinding scheme. The grids are simplicial and conforming to the fractures, in the sense that lower-dimensional cells coincide with higher-dimensional faces.
In this fully discrete formulation, the residual system (25) becomes a nonlinear system of algebraic equations, while the complementarity functions are enforced cell-wise, resulting in another system of nonlinear algebraic equations. The implementation of the generalized Newton method on the full system (or on the regularized systems in the case of IRM) is now straightforward. In GNM-RM, the return maps are performed cell-wise.
@@ 4 Numerical experiments
In this section, we will compare the performance of the three nonlinear solvers presented in section 3, GNM, IRM and GNM-RM, through a series of two- and three-dimensional numerical simulations. Specifically, we simulate pressurized fluid injection in a fractured reservoir, mimicking hydraulic stimulation of a fractured geothermal reservoir in crystalline rock. Before moving on to the experiments themselves, we will first outline the main objectives of these experiments, and provide some technical details of the implementation.
4.1 Main objectives
A central focus of the experiments is to test the impact of gradually increasing the strength of the nonlinear couplings between flow and contact mechanics on the nonlinear solver performance. Since the three nonlinear solvers under consideration are all built around the contact problem, and since GNM and IRM are already well-established for solving purely mechanical contact problems [RenardSurvey], it is important to investigate to which degree the nonlinear coupling of the contact problem to flow contributes to convergence issues. We identify the nonlinear dependence of the volumetric and hydraulic apertures on the normal displacement jump, $a_{i}=a_{\text{ref}}+[[\bm{u}]]_{n}$ and $\mathcal{A}_{i}=\mathcal{A}_{\text{ref}}+[[\bm{u}]]_{n}$, as central. Hence, we will consider three different models, with varying degrees of linearizations of these aperture laws. We will first consider a model where both the effective fracture volume and fracture permeabilities are considered constant, using $a_{i}=a_{\text{ref}}$ and $\mathcal{A}_{i}=\mathcal{A}_{\mathrm{ref}}$. Note that this only (partially) linearizes the flow equations, while the mechanical gap function (17) will still be active. Next, we consider an intermediate problem with volumetric changes, but a constant fracture permeability. This means that a constant hydraulic aperture $\mathcal{A}_{i}=\mathcal{A}_{\text{ref}}$ is used, but the volumetric aperture formula still depends on the normal displacement jump, $a_{i}=a_{\text{ref}}+[[\bm{u}]]_{n}$. Finally, we include also the cubic law on the fracture permeability and consider the fully coupled model, using the displacement-dependent formulas for both the volumetric and hydraulic apertures. These three models are summarized in table 2.
Model Feature Aperture formulas used
A Constant fracture volume and permeability $a_{i}=a_{\text{ref}},\ \mathcal{A}_{i}=\mathcal{A}_{\text{ref}}$
B Constant fracture permeability $a_{i}=a_{\text{ref}}+[[\bm{u}]]_{n}$, $\mathcal{A}_{i}=\mathcal{A}_{\text{ref}}$
C Full model $a_{i}=a_{\text{ref}}+[[\bm{u}]]_{n}$, $\mathcal{A}_{i}=\mathcal{A}_{\text{ref}}+[[\bm{u}]]_{n}$
Table 2: The three models, with varying degrees of nonlinear couplings of flow and mechanics, used in the simulations.
Another goal of the experiments is to test the sensitivity of the nonlinear solvers to the value of the augmentation parameter $c$. This is relevant to study, as the value of this parameter can greatly influence the convergence behavior of the solvers. This is demonstrated in [RenardSurvey] for purely mechanical contact problems, where GNM and IRM were shown to display sensitive behavior with respect to the value of $c$. In particular, for some of the examples, the set of $c$-values for which the solver converged was quite erratic and unpredictable.
We will consider $c$-values in a range centered around the value of 1 GPa/m. We choose $c$-values around this order, as we are aiming to balance the traction- and displacement-terms of the complementarity functions, and it is empirically found that the ratio $\bm{\lambda}/[[\bm{u}]]$ is typically in the range from $10^{-2}$ to 1 GPa/m in these simulations. This balancing of terms has been recommended in several articles on purely mechanical contact problems [hueber, gitterle]. For IRM, the value of $c$ will be gradually increased during each iteration of the outer loop; this is done because the convergence of IRM will be slower for smaller $c$-values, as mentioned in section 3.2. Specifically, if $c$ denotes the starting value, our update strategy is $c^{(k)}=\max(10^{3},10^{k}\cdot c)$, with $k$ denoting the outer loop iteration of IRM. The value is reset back to $c$ for each new time step. This feature of smaller $c$-values being slower to converge is unique to Uzawa methods like IRM, and hence, it does not apply to GNM and GNM-RM, and we simply set $c^{(k)}=c$ in these two algorithms.
Our final objective, which is related to the previous two objectives, is to assess and compare the overall performance of the three nonlinear solvers. As a related sub-objective, we will also report and discuss the performance of the linear solver, details of which can be found in the next section. The performance of each nonlinear solver will be assessed by reporting the total number of nonlinear iterations needed for convergence. By this we mean the total number of linear systems solved, summed over all time steps. For the convergence criteria, we use the same absolute tolerance of $10^{-8}$ for both the nonlinear iteration increments and residuals, both of which are measured in a Euclidean norm that is weighted by the cell volumes. We note that for IRM, the tolerance of $10^{-8}$ is used for both the outer loop and for solving the regularized system in the inner loop. An alternative would be to use a looser tolerance in the inner loop, and adaptively tighten the tolerance with increasing outer loop iterations. However, in our experiments, we did not experience any significant advantages by using a looser tolerance in the inner loop.
4.2 Technical details of the implementation
All of the simulations are run using version 1.11 of the mixed-dimensional simulation framework PorePy [porepy]. The grids are generated by Gmsh [gmsh] (version 4.13.1), while linear solver routines are imported from PETSc [balay2023petsc, DALCIN20111124] using petsc4py (version 3.24.0). Runscripts for reproducing the results of this section can be found in the Docker image in [zenodo]. The precise vertex coordinates of the fractures used in the simulations are reported in the supplementary material of this article.
The linear systems are solved by GMRES, using absolute and relative tolerances of $10^{-11}$, together with the preconditioner developed in [zabegaev2025efficient], which is tailored to mixed-dimensional poroelasticity with contact mechanics. Specifically, it is a Schur complement based preconditioner using the algebraic multigrid method, with the ILU($n$) smoother being used for the mechanics subproblem. The value of $n$ has been slightly varied across the experiments, to ensure the most robust performance of the linear solver. Specifically, we have used $n=2$ for the two-dimensional simulations and $n=0$ for the three-dimensional ones. In some cases, however, the linear solver failed to converge; in these cases, the adaptive time-stepping scheme described below is employed.
Some measures are taken to improve the conditioning of the nonlinear systems. First, to avoid having to deal with negative apertures in the non-converged state, we use a cutoff for the volumetric and hydraulic apertures that do not allow them to be less than their respective residual apertures. Thus, we are in practice using the formulas $a_{i}=\max(a_{\text{ref}},a_{\text{ref}}+[[\bm{u}]]_{n})$ and $\mathcal{A}_{i}=\max(\mathcal{A}_{\text{ref}},\mathcal{A}_{\text{ref}}+[[\bm{u}]]_{n})$. Next, we have observed empirically that the return map of GNM-RM will quite frequently result in cases where $\lVert\bm{\lambda}_{\tau}^{(k+1)}+c[[\dot{\bm{u}}^{(k+1)}]]_{\tau}\rVert=b^{(k+1)}$, i.e. the two arguments of the maximum-function of (27) are equal. In this case, GNM-RM was found to be more robust if the Jacobian corresponding to the second argument is chosen, and so this is done in all simulations. For GNM and IRM, neither of the two options showed a significant increase or decrease in the robustness of the solvers. Therefore, we arbitrarily choose the Jacobian corresponding to the second argument in case of equal arguments also for GNM and IRM. Finally, we use an appropriate scaling of the units. It is found empirically that setting the unit of mass to be $10^{9}$ kg results in the primary variables being close to the same order of magnitude, and generally improves the robustness of all solvers. Hence, for all simulations, the input data is scaled by this unit of mass.
An adaptive time-stepping scheme is used in the simulations, inspired by the one used by Simunek et al. [simunek2005hydrus]. The time-stepping scheme has two features. First, if the nonlinear solver converged, the next time step size is adapted based on the number of nonlinear iterations needed for convergence. If the number of iterations is within a specified range, the step size stays the same; if it is above this range, the step size is decreased by a factor of 0.7, while if it is below this range, the step size is increased by a factor of 3. The range will slightly vary in the examples. For the two-dimensional simulations of section 4.3, we set the range to $(4,20)$, not including the endpoints. For the three-dimensional simulations of section 4.4, we increase the lower endpoint of the range to avoid excessive computational time, setting it to $(8,20)$. In section 4.5 we run even more challenging three-dimensional simulations, in which case we shrink the range even further, to $(10,20)$.
The second feature of the time-stepping scheme is that, if the nonlinear solver fails to converge, a recomputation of the solution is attempted with a smaller step size. We use three metrics to assess non-convergent behavior: Setting the maximal number of allowed nonlinear iterations per nonlinear system to be 30, setting the maximal value of the absolute residual norm to be $10^{5}$, and setting the maximal number of allowed GMRES iterations to be 300. If any of these cases occur, the step size is decreased by a factor of one half, and a recomputation of the nonlinear system is attempted with this updated step size. If IRM fails to converge in the inner loop, the whole outer loop is restarted with the smaller time step. We allow a step size no smaller than a tenth of the initial time step, and if the simulation still fails to converge for this step size, it is terminated. We also terminate the simulation after 6 consecutive failed recomputation attempts.
Line search methods are often used as globalization strategies for nonlinear solvers. We experimented with combining the nonlinear solvers with an Armijo line search, but found that this did not improve the robustness of the solvers for the problem setups considered herein. A similar behavior has been reported for certain problems of purely mechanical contact mechanics [Acary2018]. Hence, we have chosen to run all of the simulations without a line search strategy. Finally, the material constants used in the simulations are listed in table 3, in unscaled units.
@@ 4.2 Technical details of the implementation
Residual volumetric aperture $a_{\mathrm{ref}}$ $5.0\cdot 10^{-4}$ m
Residual hydraulic aperture $\mathcal{A}_{\mathrm{ref}}$ $5.0\cdot 10^{-4}$ m
Dilation angle $\psi$ 5 degrees
@@ 4.3 Two-dimensional simulations of fluid injection
The performance of the nonlinear solvers, with the objectives of section 4.1 in mind, will first be explored in a series of two-dimensional numerical simulations. A domain of size $2000\times 1000$ m is considered, with the fracture network as shown in figure 2, and with a grid containing 33830 total cells (this includes matrix, fracture and interface cells). Gravitational effects are neglected in these two-dimensional examples. For the flow boundary conditions, we impose pressure values of 20MPa on all external boundaries. For the mechanical boundary conditions, we fix the bottom boundary (zero displacement), while compressive tractions are imposed on the remaining boundary sides. The traction values correspond to a stress tensor with non-zero values $\sigma_{xx}=-30$ MPa and $\sigma_{yy}=-50$ MPa. These pressures and anisotropic stress values are representative for crystalline rocks around 2 km depth. The initial state is chosen to be the equilibrium state, with the pressure field being uniformly equal to 20 MPa.
The pressure in the centermost cell of one of the fractures (marked by a red dot in figure 2) will momentarily increase at certain points, and otherwise be kept at a constant pressure; this mimics a pressure-controlled injection well. The injection schedule is also shown in figure 2. Starting from the background pressure of 20 MPa, the overpressure in the well is increased by 1 MPa every hour, performing the first pressure increase at $t_{\text{start}}=0$, and running the simulation until $t_{\text{end}}=3$h. We will refer to the period between each pressure increase as the different phases of the simulation. To accurately capture the dynamics immediately after each pressure increase, a small time step of one second is initially used. The adaptive time-stepping scheme described earlier is subsequently employed.
Figure 2: Left: Fracture network used in the two-dimensional simulations, with the injection well marked by a red dot. Right: Injection schedule for the two-dimensional simulations. The wellhead pressure is momentarily increased by 1 MPa every hour, and otherwise kept constant. The same injection schedule will also be used for the three-dimensional simulations in section 4.4
We run simulations with the setup described above, using all three models of table 2. Figure 3 provides a visualization of the simulation at the end of each phase of the injection schedule using model C. For comparison, we also include in the figure the end of the final phase using models A and B. The plots show the pressure profile and contact states (left column), as well as the changes in volumetric aperture and displacement magnitudes relative to the initial state at $t=0$ (right column). We observe that each pressure increase in the well causes a larger portion of the fractures to slip. In the cases where models B and C are used, the slip events cause an increase in the volumetric aperture (this obviously does not happen in model A, where the volumetric aperture is constant). The increase in volumetric aperture in models B and C occurs due to the shear dilation of fractures causing an increase in the normal displacement jump. All models result in mostly the same fracture cells slipping, though with some minor differences. We also note that the pressures, displacement and volumetric apertures all gradually increase from models A through C.
In figure 4, we report on the cumulative number of linear and nonlinear iterations, summed over all time steps, for the different nonlinear solvers and for different values of the augmentation parameter. As explained earlier, an update strategy is used on the augmentation parameter for IRM, and so the $c$-values in figure 4 refer to the starting values.
Figure 3: Visualization of the simulations of two-dimensional injection. Results are shown at the end of each simulation phase using model C, and at the end of the final simulation phase using models A and B. Left: Pressure profile and contact states. Here, slip is measured in a cumulative manner; a cell is regarded as being in a ”slip” state if the fracture is closed and the norm of the tangential displacement jump is positive, using the state before the injection (at $t=0$) as the reference. There is no opening of fractures. Right: Changes in volumetric aperture and displacement magnitude, relative to the respective values at $t=0$
Figure 4: Bar plots showing the cumulative number of nonlinear iterations, summed over all time steps, for the example of section 4.3 with the different degrees of nonlinear couplings between flow and mechanics outlined in table 2. The three different colors per bar correspond to the three different phases of the simulation, after each pressure increase in the injection well. Hatched black lines indicate steps where the nonlinear solver failed to converge, and red crosses indicate a simulation that was terminated. The simulation is terminated if one of three cases occur: The number of nonlinear iterations exceed 800, the nonlinear solver failed to converge for the smallest allowed time step, or after 6 unsuccessful recomputation attempts. Finally, the total number of iterations of the linear solver is marked by a black dot
With the objectives of section 4.1 in mind, we start by discussing the impact of the different constitutive laws for the apertures. From figure 4, we can clearly observe an increase in difficulty from model A through model C, with the number of iterations needed for convergence, as well as the number of non-convergent cases, generally increasing (note that the simulation is terminated if the number of nonlinear iterations exceeds 800; these cases may have converged if more iterations were used, but they are deemed too slow). The increase in difficulty is most pronounced between models A and B, with a sharp increase in the number of cases where the nonlinear solvers do not converge. We note in particular that IRM performs very poorly on models B and C, with no cases converging within 800 iterations, and several cases not even getting past the first injection phase. Overall, these results indicate that the nonlinear coupling between flow and mechanics, induced by the displacement-dependent aperture formulas, significantly adds to the difficulty of the problem, and that IRM is worse at handling this nonlinear coupling than the other two solvers.
The performance of the linear solver is discussed next. We note from figure 4 a significant increase in the average number of GMRES iterations per nonlinear iteration from models A to B. This is consistent with the observation that the overall difficulty of the problem is significantly increased between these two models. Another interesting observation is that for IRM, the average number of GMRES iterations per nonlinear iteration generally decreases with increasing $c$-value, while it is roughly constant for GNM and GNM-RM ($c=10$ is an exception for IRM; here the number slightly increases again). Finally, it is worth noting that for models B and C, GNM-RM uses consistently less GMRES iterations per nonlinear iteration than GNM. Hence, for the cases considered here, the return map results in a better performance of the preconditioned linear solver.
Next, we discuss the impact of the augmentation parameter on the solver performances. First, we observe from figure 4(a) the feature of IRM being slower to converge for lower $c$-values, which is present even with our relatively aggressive update strategy on the $c$-value. This feature, combined with the fact that IRM must solve more nonlinear systems per time step, causes IRM to be noticeably slower than GNM and GNM-RM. Second, we observe that both GNM and IRM perform poorly for the largest $c$-values in most cases, where they are either not converging at all, or using an excessive amount of nonlinear iterations. Additionally, IRM fails to converge also for the two smallest $c$-values for models B and C. GNM-RM, on the other hand, converges for all $c$-values, and is also quite consistent across the range of $c$-values, in regards to the number of nonlinear iterations. These results indicates that GNM-RM is less sensitive to the value of the augmentation parameter than the other two solvers, and in particular, that it handles higher values of the parameter much better.
A closer inspection of the cases where GNM and IRM fail to converge (hatched lines in figure 4) reveals a correlation between the mode of failure and the value of the augmentation parameter. For the higher values ($c>0$), the failure mode is nearly always cycling of the iterations, while for the lower values ($c<0$), the iterations either diverge to infinity or simply stagnate with no apparent pattern. Examples of both cases are shown in figure 5, in which residual norms are plotted for the simulation using model A and a $c=10^{3}$, and the simulation using model C and $c=10^{-3}$. Different modes of failure are observed in the two cases. In the case with the higher $c$-value, GNM and IRM fails by cycling of the iterations, illustrated by repetitive patterns in the residual history. In the case with the lower $c$-value, IRM instead fails by the iterations diverging to infinity. GNM-RM is seen to converge quickly in both cases. An interesting conclusion to be drawn here is that, for these examples, performing a return map after each generalized Newton iteration is an effective way of avoiding cycling of the iterations.
Figure 5: Residual norms at every nonlinear iteration for the simulation using model A and $c=10^{3}$ (left), and the simulation using model C and $c=10^{-3}$ (right). Only the first attempted time step is shown for all solvers
In light of the entire comparison, we conclude that for these examples, IRM is the worst-performing solver, while GNM-RM is the best-performing one. IRM is both slower to converge than the other two solvers, as well as having more convergence issues for the models with the stronger nonlinear couplings between flow and mechanics. In terms of the number of nonlinear iterations, GNM and GNM-RM perform relatively similar. GNM-RM has the added benefit of being less sensitive to the value of the augmentation parameter than GNM.
The scalability of the solvers with respect to the grid resolution is also of interest. We performed an experiment testing the effect of the grid resolution on the number of nonlinear iterations needed for convergence. We have chosen not to include this experiment, however, as we were not able to draw any reliable conclusions from it. The general tendency was a weak grid dependency, with a modest increase in the number of iterations for finer grids in most cases. However, this pattern was also broken on several occasions, as the grid refinement would sometimes result in a fewer number of iterations.